OpenStudy (anonymous):

Help !? The wavelengths of visible light are between m and m. Are these wavelengths longer or shorter than those of ultraviolet rays? Explain. (Hint: Make sure that your answer from part a is in scientific notation when comparing the values.) 2 POINTS

OpenStudy (jamesj):

between what and what?

OpenStudy (jamesj):

...around 390 - 750 nanometers.

OpenStudy (jamesj):

A nano-meter is \( 10^{-9} \) meters. Now, if UV light is around \( 10^{-8} - 10^{-9} \) meters, is visible light of longer or shorter wavelength?

OpenStudy (anonymous):

Oh My Bad ! its 3.8x10^-7 m And 7.6x10^-7

OpenStudy (jamesj):

Right. Hence that's longer or shorter than UV light?

OpenStudy (anonymous):

don know , thats what im trying 2 figure out lol

OpenStudy (jamesj):

\[ ... > 10^3 > 10^2 > 10^1 > 10^0 = 1 > 10^{-1} = \frac{1}{10} > 10^{-2} > ... 10^{-7} > 10^{-8} > 10^{-9} > ... \]

OpenStudy (jamesj):

So is the wavelength of visible light, for example \[ 3.8x10^{-7} \ m \] is that number larger or smaller than the wavelength of UV light, for example, \[ 10^{-8} \ m \] ?

OpenStudy (jamesj):

\[ 3.8 \times 10^{-7} \ m \]

OpenStudy (jamesj):

Remember: \[ 10^{-7} > 10^{-8} \]

OpenStudy (anonymous):

Uhhh , Smaller ? I Dont Get This -_-

OpenStudy (jamesj):

...and that's because \[ 10^{-8} = 10^{-1} \times 10^{-7} = \frac{1}{10} \times 10^{-7} < 10^{-7} \]

OpenStudy (jamesj):

Ok. Let's take a step back. Which is larger: 10^2 or 10^1 ?

OpenStudy (anonymous):

10^2 .

OpenStudy (jamesj):

right. 10^2 is larger because 10^2 = 100, but 10^1 = 10 and 100 > 10

OpenStudy (jamesj):

Now which is larger: 1 or 10^-1 ?

OpenStudy (anonymous):

Ok :D Got Tht.

OpenStudy (anonymous):

10^-1

OpenStudy (jamesj):

No. \[ 10^{-1} = \frac{1}{10} \] and 1/10 is less than 1.

OpenStudy (jamesj):

make sense?

OpenStudy (anonymous):

Ohhh lol Ok

OpenStudy (jamesj):

Now, which is larger: \[ 10^{-10} \] or \[ 10^{-12} \] ?

OpenStudy (anonymous):

10^-10?

OpenStudy (jamesj):

Yes, because \[ 10^{-12} = 10^{-2} \times 10^{-10} \] and \[ 10^{-2} = \frac{1]{100} \] Therefore \[ 10^{-12} < 10^{-10} \]

OpenStudy (jamesj):

\[ 10^{-2} = \frac{1}{100} \]

OpenStudy (anonymous):

Ohh , Ok

OpenStudy (jamesj):

Now, if UV light has a wavelength of around \( 10^{-8} \) meters, and visible light has a wavelength around \( 10^{-7} \), which has the longer wavelength?

OpenStudy (anonymous):

10^-7 ?

OpenStudy (jamesj):

yes, visible light.

OpenStudy (anonymous):

:D

OpenStudy (jamesj):

Negative exponents are a bit tricky, but just remember the negative means reciprocal. Hence something like this is true: \[ 10^{-1} > 10^{-123748123704} \]

OpenStudy (anonymous):

Ohhhhh Yea

OpenStudy (jamesj):

Hence if \[ 10^p > 10^q \] then \[ 10^{-p} < 10^{-q} \]

OpenStudy (jamesj):

e.g., \[ 10^7 > 10^2 \] but \[ 10^{-7} < 10^{-2} \] because \[ \frac{1}{10^7} < \frac{1}{10^2} \]

OpenStudy (anonymous):

Yea i Get it , So Is My Answer Longer Or Shorter?

OpenStudy (anonymous):

:D

OpenStudy (jamesj):

10^-7 meters is longer or shorter than 10^-8 meters ?

OpenStudy (anonymous):

ohh , longer lol.

OpenStudy (jamesj):

Longer, yes.

OpenStudy (anonymous):

Thnks :D